this is the word problem. juliet has a choice between recieving an $1800 monthly salary or a base salary of $1600 plus a 4% commission on the sales she has during the month. for what amount of sales will the two choices be equal?
Answer by mananth(16946)

Salary = 1800
commission+ salary = 4%=0.04 +1600
sales = $x
salary = salary + commision
1800= 1600+0.04x
200=0.04x
x=200/0.04
x= $5000 sales